Frost wedging breaks apart a large rock measuring 12 m long by 12 m wide by 2 m high into four equal pieces measuring 6 m long by 6 m wide by 2 m high. How much greater, in m2, is the total surface area of the four pieces compared to that of the original rock?
Answer #1

The surface area of the bigger piece of rock = 2(12*12)+ 2(12*2)+ 2(2*12) = 384 m2

The surface area of one of the smaller pieces = 2(6*6)+ 2(6*2)+ 2(2*6) = 120m2

The surface area of four such pieces = 4*120 =480m2

The total surface area of the four pieces compared to that of the original rock is 1.25 times.